# How many measurements are enough? Bayesian recovery in inverse problems with general distributions

### v1

@@ 1 Introduction
Here $e\in\mathbb{R}^{m}$ is measurement noise and $A\in\mathbb{R}^{m\times n}$, often termed the measurement matrix, represents the forwards operator. While simple, the discrete, linear problem (1.1) is sufficient to model many important applications burger2024learning ; scarlett2022theoretical ; ongie2020deep . It is common to solve (1.1) using a Bayesian approach (see, e.g., dashti2017bayesian ; stuart2010inverse ). Here one assumes that $x^{*}$ is drawn from some prior distribution $\mathcal{R}$. However, in practice, $\mathcal{R}$ is never known exactly. Especially in modern settings that employ Deep Learning (DL) arridge2019solving ; dimakis2022deep , it is common to learn an approximate prior $\mathcal{P}$ and then recover $x^{*}$ from $y$ by approximate posterior sampling, i.e., sampling $\hat{x}$ from the posterior $\mathcal{P}(\cdot|y,A)$. An increasingly popular approach involves using generative models to learn $\mathcal{P}$ (see, e.g., arridge2019solving ; dimakis2022deep ; scarlett2022theoretical ; zhao2023generative ; bohra2023bayesian ; song2022solving and references therein).
@@ 1.1 Overview
In this work, we strive to answer to this question in the broadest possible terms, with a theoretical framework that allows for very general types of distributions. We now describe the corresponding conditions needed and present simplified a version of our main results.
(i) Closeness of the real and approximate distributions. Namely, $W_{p}(\mathcal{R},\mathcal{P})$ is small, where $W_{p}$ denotes the Wassertstein $p$-metric for some $1\leq p\leq\infty$.
@@
for all $x\in D:=\mathrm{supp}(\mathcal{P})-\mathrm{supp}(\mathcal{P})$ and $t>0$. If $\mathcal{A}$ is isotropic, i.e., $\mathbb{E}_{A\sim\mathcal{A}}{\|Ax\|}^{2}={\|x\|}^{2}$, $\forall x\in\mathbb{R}^{n}$, as is often the case in practice, these constants measure how fast ${\|Ax\|}^{2}$ concentrates around its mean. Notice that this condition is imposed only on $D=\mathrm{supp}(\mathcal{P})-\mathrm{supp}(\mathcal{P})$, rather than the whole space $\mathbb{R}^{n}$. As we see later, this is crucial in obtaining meaningful recovery guarantees.
@@
(iv) Gaussian noise. Namely, $\mathcal{E}=\mathcal{N}(0,\frac{\sigma^{2}}{m}I)$ where $\sigma>0$.
@@
Let $1\leq p\leq\infty$, $40<\delta\leq 1/$ 4, $\eta>0$ and suppose that conditions (i)–(v) hold with $W_{p}(\mathcal{R},\mathcal{P})\leq\varepsilon/(2m\theta)$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Suppose that $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ independently and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$, where $y=Ax^{*}+e$. Then, for any $d\geq 2$,
@@
Note that in this and all subsequent results, the term of the left-hand side of (1.2) is the probability with respect to all variables, i.e., $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$. This result is extremely general, in that it allows for essentially arbitrary (real and approximate) signal distributions $\mathcal{R}$ and $\mathcal{P}$ and an essentially arbitrary distribution $\mathcal{A}$ for the forwards operator. In broad terms, it bounds the probability that the error ${\|x^{*}-\hat{x}\|}$ of posterior sampling exceeds a constant times the noise level $\sigma$ plus an arbitrary parameter $\eta$. It does so in terms of the approximate covering number $\mathrm{Cov}_{\eta,\delta}(\mathcal{P})$, which measures the complexity of the approximate distribution $\mathcal{P}$, the concentration bounds $C_{\mathsf{low}}$ and $C_{\mathsf{upp}}$ for $\mathcal{A}$, which measure how much $A$ elongates or shrinks a fixed vector, and an exponentially-decaying term $\mathrm{e}^{-m/16}$, which stems from the (Gaussian) noise. In particular, by analyzing these terms for different classes of distributions $\mathcal{P}$ and $\mathcal{A}$, we can derive concrete bounds for various exemplar problems. We next describe two such problems.
@@ 1.2 Examples
Later in Theorem 3.5, we slightly refine and generalize this result. This theorem shows the efficacy of Bayesian recovery with subgaussian random matrices: namely, the sample complexity scales linearly in the distribution complexity, i.e., the log of the approximate covering number.
@@
Consider the setup of Theorem 1.1, where $\mathcal{A}$ is a distribution of subsampled orthogonal matrices based on matrix $U$. Then there is a universal constant $c>0$ such that
@@
See Theorem 3.8 for the full result. This result shows that similar stable and accurate recovery to the Gaussian case can be achieve using subsampled orthogonal matrices, provided the number of measurements scales like the coherence. We discuss this term further in §3.2.
Theorems 1.2 and 1.3 establish a key relationship for successful Bayesian recovery in inverse problems, in each case relating the number of measurements $m$ to the intrinsic complexity $\log(\mathrm{Cov}_{\eta,\delta}(\mathcal{P}))$ of $\mathcal{P}$. It is therefore informative to see how this complexity behaves for cases of interest.
As noted, it is common to use a generative model to learn $\mathcal{P}$. This means that $\mathcal{P}=G\sharp\gamma$, where $G:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is a Deep Neural Network (DNN) and $\gamma$ is some fixed probability measure on the latent space $\mathbb{R}^{k}$. Typically, $\gamma=\mathcal{N}(0,I)$. If $G$ is $L$-Lipschitz, we show in Proposition 4.1 that
@@
This scales log-linearly in the latent space dimension $k$, confirming the intrinsic low-complexity of the distribution $\mathcal{P}$. Combining (1.3) with Theorems 1.2 and 1.3 we see that posterior sampling achieves stable and accurate recovery, provided the number of measurements $m$ scales log-linearly (and thereby near-optimally) with $k$.
@@ 1.3 Significance
1. We provide the first results for Bayesian recovery with arbitrary distributions real and approximate prior distributions $\mathcal{R}$, $\mathcal{P}$ and forwards operator and noise distributions $\mathcal{A}$, $\mathcal{E}$.
2. Unlike much of the theory of Bayesian inverse problems, which is asymptotic in nature arridge2019solving ; dashti2017bayesian ; stuart2010inverse , our results are non-asymptotic. They hold for arbitrary values of the various paremeters, within given ranges (e.g., the failure probability $\delta$, the noise level $\sigma$, the number of measurements $m$, and so forth).
@@
4. For the important class of subsampled orthogonal transforms, we show the sample complexity of Bayesian recovery depends on the coherence, thus resolving several key open problems in the literature (see next).
5. It is increasingly well-known that DL-based methods for inverse problems are susceptible to hallucinations and other undesirable effects reader2023ai ; hoffman2021promise ; belthangady2019applications ; bhadra2021hallucinations ; antun2020instabilities ; colbrook2022difficulty ; gottschling2025troublesome ; burger2024learning ; genzel2023solving ; huang2018some ; morshuis2022adversarial ; raj2020improving ; zhang2021instabilities ; noordman2023complexities ; muckley2021results . This is a major issue that may limit the uptake of these methods in safety-critical domains such as medical imaging burger2024learning ; muckley2021results ; laine2021avoiding ; liu2022medical ; varoquaux2022machine ; wu2021medical ; yu2022validation . Our results provide theoretical guarantees for stable and accurate, and therefore show conditions under which hallucinations provably cannot occur. This is not only theoretically interesting, but it also has practical consequences in the development of robust DL methods for inverse problems – a topic we intend to explore in future work.
@@ 1.4 Related work
Bayesian methods for inverse problems have become increasingly popular over the last several decades dashti2017bayesian ; stuart2010inverse . Many state-of-the-art DL methods for inverse problems now follow a Bayesian approach (see arridge2019solving ; adler2025deep ; laumont2022bayesian ; holden2022bayesian ; bohra2023bayesian ; ongie2020deep and references therein). Moreover, learned priors such as those stemming from generative models, are now increasingly used in applications arridge2019solving ; dimakis2022deep ; jalal2021robust ; song2022solving ; adler2025deep ; holden2022bayesian ; bohra2023bayesian ; kadkhodaie2021stochastic ; aali2023solving ; kawar2022denoising ; chung2022score ; feng2023score ; luo2023bayesian ; zach2023stable ; zhao2023generative ; kawar2021snips ; scarlett2022theoretical ; ongie2020deep .
@@
This work is motivated in part by the (non-Bayesian) theory of generative models for solving inverse problems (see bora2017compressed ; shah2018solving ; hand2018phase ; hand2018global ; dimakis2022deep ; jalal2021robust ; song2022solving ; zhao2023generative ; scarlett2022theoretical and references therein). This was first developed in bora2017compressed , where compressed sensing techniques were used to show recovery guarantees for a Gaussian random matrix $A$ when computing an approximate solution to (1.1) in the range $\Sigma:=\mathrm{ran}(G)$ of a Lipschitz map $G:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$, typically assumed to be a generative NN. This is a deterministic approach. Besides the random forward operator and (potentially) random noise, it recovers a fixed (i.e., nonrandom) underlying signal $x^{*}$ in a deterministic fashion with a point estimator $\hat{x}$ that is obtained as a minimizer of the empirical $\ell^{2}$-loss $\min_{z\in\Sigma}{\|Az-y\|}^{2}$. In particular, no information about the latent space distribution $\gamma$ is used. In this work, following, e.g., jalal2021instance ; jalal2021robust ; song2022solving ; bohra2023bayesian ; aali2023solving and others, we consider a Bayesian setting, where $x^{*}\sim\mathcal{R}$ is random and where we quantify the number of measurements that suffice for accurate and stable recovery via posterior sampling $\hat{x}\sim\mathcal{P}(\cdot|y,A)$.
@@
Our work is a generalization of jalal2021instance , which considered Bayesian recovery in with Gaussian random matrices and standard Gaussian noise. We significantly extend their results to allow for arbitrary distributions $\mathcal{A}$ and $\mathcal{E}$ for the forward operator and noise, respectively. Using Theorem 1.1, we derive guarantees for the case of subsampled orthogonal transforms (Theorem 1.3, which, unlike Gaussian random matrices, is very relevant to applications. In particular, if $U$ is a DFT matrix our work addresses open problems posed in (jalal2021robust, , §3) and (scarlett2022theoretical, , §II.F) on recovery guarantees with Fourier measurements. Moreover, addressing an open problem posed in (jalal2021instance, , §6), we derive concrete bounds for the approximate covering number of distributions given by Lipschitz generative DNNs (see (1.3) and Proposition 4.1) and distributions of sparse vectors (see (1.4) and Proposition 4.3). In particular, we demonstrate stable and accurate recovery in a Bayesian sense with a number of measurements that scales linearly in the model complexity, i.e., $k$ in the former case and $s$ in the latter case.
@@
Recently berk2022coherence extended the results of bora2017compressed in the non-Bayesian setting from Gaussian random matrices to subsampled orthogonal transforms when considering ReLU generative DNNs. Theorem 1.3 provides a Bayesian analogue of this work. We also generalize berk2022coherence by allowing for general measurement distributions $\mathcal{A}$ and priors $\mathcal{P}$, which need not be arise from ReLU generative DNNs.
@@
Classical compressed sensing considers the recovery of approximately $s$-sparse vectors from (1.1). However, it has been extended to consider much more general types of low-complexity signal models, such as joint or block sparse vectors, tree sparse vectors, cosparse vectors and many others adcock2017breaking ; baraniuk2010model-based ; bourrier2014fundamental ; davenport2012introduction ; duarte2011structured ; traonmilin2018stable . However, most recovery guarantees for general model classes consider only (sub)Gaussian random measurements (see e.g., baraniuk2010model-based ; dirksen2016dimensionality ). Recently, adcock2024unified introduced a general framework for compressed sensing that allows for essentially arbitrary low-complexity models and arbitrary (random) measurement matrices. Our is a Bayesian analogue of this deterministic work. Similar to adcock2024unified , a key feature of our work is that we consider arbitrary real and approximate signal distributions $\mathcal{P},\mathcal{R}$ (analogous to arbitrary low-complexity models) and arbitrary distributions $\mathcal{A}$ for the forwards operator. Unsurprisingly, a number of the conditions for our main result, Theorem 1.1 – namely, the low-complexity condition (ii) and concentration condition (iii) – share similarities to those that ensure stable and accurate recovery in non-Bayesian compressed sensing. See Remarks 2.2 and 3 for further discussion. However, the proof techniques used in this work are entirely different.

### v2

@@ 1 Introduction
Here $e\in\mathbb{R}^{m}$ is measurement noise and $A\in\mathbb{R}^{m\times n}$, often termed the measurement matrix, represents the forwards operator. While simple, the discrete, linear problem (1.1) is sufficient to model many important applications adcock2021compressive ; burger2024learning ; scarlett2022theoretical ; ongie2020deep . It is common to solve (1.1) using a Bayesian approach (see, e.g., dashti2017bayesian ; stuart2010inverse ), where one assumes that $x^{*}$ is drawn from some prior distribution $\mathcal{R}$. However, in practice, $\mathcal{R}$ is never known exactly. Especially in modern settings that employ Deep Learning (DL) arridge2019solving ; dimakis2022deep , it is typical to learn an approximate prior $\mathcal{P}$ and then recover $x^{*}$ from $y$ by approximate posterior sampling, i.e., sampling $\hat{x}$ from the posterior $\mathcal{P}(\cdot|y,A)$. An increasingly popular approach involves using generative models to learn $\mathcal{P}$ (see, e.g., arridge2019solving ; dimakis2022deep ; scarlett2022theoretical ; zhao2023generative ; bohra2023bayesian ; song2022solving and references therein).
@@ 1.1 Overview
In this work, we strive to answer to this question in the broadest possible terms, with a theoretical framework that allows for very general types of distributions. We now describe the corresponding conditions needed and present simplified versions of our main results.
(i) Closeness of the real and approximate distributions. We assume that $W_{p}(\mathcal{R},\mathcal{P})$ is small for some $1\leq p\leq\infty$, where $W_{p}$ denotes the Wassertstein $p$-metric.
@@
for all $x\in D:=\mathrm{supp}(\mathcal{P})-\mathrm{supp}(\mathcal{P})$ and $t>0$. Here, and throughout this paper, we write $S-S=\{x_{1}-x_{2}:x_{1},x_{2}\in S\}\subseteq\mathbb{R}^{n}$ for the difference set associated with a set $S\subseteq\mathbb{R}^{n}$. We also write $\mathrm{supp}(\mathcal{P})$ for the support of the measure $\mathcal{P}$ (see §2.1 for the definition). Furthermore, we write $\left\|\cdot\right\|$} for the $\ell^{2}$-norm and $\left\|\cdot\right\|$$_{\infty}$ for the $\ell^{\infty}$-norm. If $\mathcal{A}$ is isotropic, i.e., $\mathbb{E}_{A\sim\mathcal{A}}{\|Ax\|}^{2}={\|x\|}^{2}$, $\forall x\in\mathbb{R}^{n}$, as is often the case in practice, these constants measure how fast ${\|Ax\|}^{2}$ concentrates around its mean. Notice that this condition is imposed only on $D=\mathrm{supp}(\mathcal{P})-\mathrm{supp}(\mathcal{P})$, rather than the whole space $\mathbb{R}^{n}$. As we see later, this is crucial in obtaining meaningful recovery guarantees.
@@
(iv) Gaussian noise. Specifically, we assume that $\mathcal{E}=\mathcal{N}(0,\frac{\sigma^{2}}{m}I)$ for some $\sigma>0$.
@@
Let $1\leq p\leq\infty$, $0<\delta\leq 1/4$, $\varepsilon,\eta>0$ and suppose that conditions (i)–(v) hold with $W_{p}(\mathcal{R},\mathcal{P})\leq\varepsilon/(2m\theta)$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Suppose that $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ independently and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$, where $y=Ax^{*}+e$. Then, for any $d\geq 2$,
@@
Note that in this and all subsequent results, the term on the left-hand side of (1.2) is the probability with respect to all variables, i.e., $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$. Theorem 1.1 is extremely general, in that it allows for essentially arbitrary (real and approximate) signal distributions $\mathcal{R}$ and $\mathcal{P}$ and an essentially arbitrary distribution $\mathcal{A}$ for the forwards operator. In broad terms, it bounds the probability that the error $^{\|x^{*}-\hat{x}\|}$ of posterior sampling exceeds a constant times the noise level $\sigma$ plus an arbitrary parameter $\eta$. It does so in terms of the approximate covering number $\mathrm{Cov}_{\eta,\delta}(\mathcal{P})$, which measures the complexity of the approximate distribution $\mathcal{P}$, the concentration bounds $C_{\mathsf{low}}$ and $C_{\mathsf{upp}}$ for $\mathcal{A}$, which measure how much $A$ elongates or shrinks a fixed vector, and an exponentially-decaying term $\mathrm{e}^{-m/16}$, which stems from the (Gaussian) noise. In particular, by analyzing these terms for different classes of distributions $\mathcal{P}$ and $\mathcal{A}$, we can derive concrete bounds for various exemplar problems. We next describe two such problems.
@@ 1.2 Examples
This theorem shows the efficacy of Bayesian recovery with subgaussian random matrices: namely, the sample complexity scales linearly in the distribution complexity, i.e., the log of the approximate covering number. Later in Theorem 3.5, we slightly refine and generalize this result.
@@
Consider the setup of Theorem 1.1, where $\mathcal{A}$ is a distribution of subsampled orthogonal matrices based on a matrix $U$. Then there is a universal constant $c>0$ such that
@@
This result shows that similar stable and accurate recovery to the Gaussian case can be achieved using subsampled orthogonal matrices, provided the number of measurements scales linearly with the coherence. We discuss this term further in §3.2, where we also present the full result, Theorem 3.9.
In general, Theorems 1.2 and 1.3 establish a key condition for successful Bayesian recovery in inverse problems, in each case relating the number of measurements $m$ to the intrinsic complexity $\log(\mathrm{Cov}_{\eta,\delta}(\mathcal{P}))$ of $\mathcal{P}$. It is therefore informative to see how this complexity behaves for cases of interest. As noted, it is common to use a generative model to learn $\mathcal{P}$. This means that $\mathcal{P}=G\sharp\gamma$, where $G:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is a Deep Neural Network (DNN) and $\gamma$ is some fixed probability measure on the latent space $\mathbb{R}^{k}$. Typically, $\gamma=\mathcal{N}(0,I)$. If $G$ is $L$-Lipschitz, we show in Proposition 4.1 that
@@
This scales log-linearly in the latent space dimension $k$, confirming the intrinsic low-complexity of $\mathcal{P}$. Combining (1.3) with Theorems 1.2-1.3 we see that posterior sampling achieves stable and accurate recovery, provided the number of measurements $m$ scales near-optimally with $k$, i.e., as $\mathcal{O}(k\log(k))$.
@@ 1.3 Significance
1. We provide the first results for Bayesian recovery with arbitrary real and approximate prior distributions $\mathcal{R}$, $\mathcal{P}$ and forwards operator and noise distributions $\mathcal{A}$, $\mathcal{E}$.
2. Unlike much of the theory of Bayesian inverse problems, which is asymptotic in nature arridge2019solving ; dashti2017bayesian ; stuart2010inverse , our results are non-asymptotic. They hold for arbitrary values of the various parameters within given ranges (e.g., the failure probability $\delta$, the noise level $\sigma$, the number of measurements $m$, and so forth).
@@
4. For the important class of subsampled orthogonal transforms, we show that the sample complexity of Bayesian recovery depends on the coherence, thus resolving several key open problems in the literature (see next).
5. It is increasingly well-known that DL-based methods for inverse problems are susceptible to hallucinations and other undesirable effects neyra-nesterenko2023stable ; reader2023ai ; hoffman2021promise ; belthangady2019applications ; bhadra2021hallucinations ; antun2020instabilities ; colbrook2022difficulty ; gottschling2025troublesome ; burger2024learning ; genzel2023solving ; huang2018some ; morshuis2022adversarial ; raj2020improving ; zhang2021instabilities ; noordman2023complexities ; muckley2021results . This is a major issue that may limit the uptake of these methods in safety-critical domains such as medical imaging burger2024learning ; muckley2021results ; laine2021avoiding ; liu2022medical ; varoquaux2022machine ; wu2021medical ; yu2022validation . Our results provide theoretical guarantees for stable and accurate, and therefore show conditions under which hallucinations provably cannot occur. This is not only theoretically interesting, but it also has practical consequences in the development of robust DL methods for inverse problems – a topic we intend to explore in future work.
@@ 1.4 Related work
Bayesian methods for inverse problems have become increasingly popular over the last several decades dashti2017bayesian ; stuart2010inverse , and many state-of-the-art DL methods for inverse problems now follow a Bayesian approach (see arridge2019solving ; adler2025deep ; laumont2022bayesian ; holden2022bayesian ; bohra2023bayesian ; ongie2020deep and references therein). Learned priors, such as those stemming from generative models, are now also increasingly used in applications arridge2019solving ; dimakis2022deep ; jalal2021robust ; song2022solving ; adler2025deep ; holden2022bayesian ; bohra2023bayesian ; kadkhodaie2021stochastic ; aali2023solving ; kawar2022denoising ; chung2022score ; feng2023score ; luo2023bayesian ; zach2023stable ; zhao2023generative ; kawar2021snips ; scarlett2022theoretical ; ongie2020deep .
@@
This work is motivated in part by the (non-Bayesian) theory of generative models for solving inverse problems (see bora2017compressed ; shah2018solving ; hand2018phase ; hand2018global ; dimakis2022deep ; jalal2021robust ; song2022solving ; zhao2023generative ; scarlett2022theoretical and references therein). This was first developed in bora2017compressed , where compressed sensing techniques were used to show recovery guarantees for a Gaussian random matrix $A$ when computing an approximate solution to (1.1) in the range $\Sigma:=\mathrm{ran}(G)$ of a Lipschitz map $G:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$, typically assumed to be a generative DNN. This is a deterministic approach. Besides the random forward operator and (potentially) random noise, it recovers a fixed (i.e., nonrandom) underlying signal $x^{*}$ in a deterministic fashion with a point estimator $\hat{x}$ that is obtained as a minimizer of the empirical $\ell^{2}$-loss $\min_{z\in\Sigma}{\|Az-y\|}^{2}$. In particular, no information about the latent space distribution $\gamma$ is used. In this work, following, e.g., jalal2021instance ; jalal2021robust ; song2022solving ; bohra2023bayesian ; aali2023solving and others, we consider a Bayesian setting, where $x^{*}\sim\mathcal{R}$ is random and where we quantify the number of measurements that suffice for accurate and stable recovery via posterior sampling $\hat{x}\sim\mathcal{P}(\cdot|y,A)$.
@@
Our work is a generalization of jalal2021instance , which considered Bayesian recovery with Gaussian random matrices and standard Gaussian noise. We significantly extend jalal2021instance to allow for arbitrary distributions $\mathcal{A}$ and $\mathcal{E}$ for the forward operator and noise, respectively. A key technical step in doing this is the introduction of the concentration bounds $C_{\mathsf{low}}$ and $C_{\mathsf{upp}}$. In particular, these bounds are imposed only over the subset $D=\mathrm{supp}(\mathcal{P})-\mathrm{supp}(\mathcal{P})$. This is unnecessary in the Gaussian case considered in Theorem 1.2, but crucial to obtain meaningful bounds in, for instance, the case of subsampled orthogonal transforms considered in Theorem 1.3 (see Remarks 3 and 3.2 for further discussion). As noted, this case is very relevant to applications. In particular, when $U$ is taken as a DFT matrix our work addresses open problems posed in (jalal2021robust, , §3) and (scarlett2022theoretical, , §II.F) on recovery guarantees with Fourier measurements. We also derive bounds for the approximate covering number of distributions given by Lipschitz generative DNNs (see (1.3) and Proposition 4.1) and distributions of sparse vectors (see (1.4) and Proposition 4.3), addressing an open problem posed in (jalal2021instance, , §6). In particular, we demonstrate stable and accurate recovery in a Bayesian sense with a number of measurements that scales linearly in the model complexity, i.e., $k$ in the former case and $s$ in the latter case.
@@
Recently berk2022coherence ; berk2023model generalized the results of bora2017compressed in the non-Bayesian setting from Gaussian random matrices to subsampled orthogonal transforms. Theorem 1.3 provides a Bayesian analogue of this work, as discussed above, where we consider posterior sampling rather than a deterministic point estimator. We also extend the setup ofberk2022coherence ; berk2023model by allowing for general measurement distributions $\mathcal{A}$ and priors $\mathcal{P}$. In particular, in berk2022coherence ; berk2023model the quantity $\Sigma$, which is the deterministic analogue of the prior distribution $\mathcal{P}$ in our work, was assumed to be the range of a ReLU generative DNN. Like in berk2022coherence , we make use of the concept of coherence (see Theorem 1.3). However, our proof techniques are completely different to those of berk2022coherence ; berk2023model .
@@
Classical compressed sensing considers the recovery of approximately $s$-sparse vectors from (1.1). However, it has been extended to consider much more general types of low-complexity signal models, such as joint or block sparse vectors, tree sparse vectors, cosparse vectors and many others adcock2019joint ; adcock2017breaking ; baraniuk2010model-based ; bourrier2014fundamental ; davenport2012introduction ; duarte2011structured ; traonmilin2018stable . However, most recovery guarantees for general model classes consider only (sub)Gaussian measurement matrices (see e.g., baraniuk2010model-based ; dirksen2016dimensionality ). Recently, adcock2024unified introduced a general framework for compressed sensing that allows for general low-complexity models $\Sigma\subseteq\mathbb{R}^{n}$ contained within a finite union of finite-dimensional subspaces and arbitrary (random) measurement matrices. Our work is a Bayesian analogue of this deterministic setting. Similar to adcock2024unified , a key feature of our work is that we consider arbitrary real and approximate signal distributions $\mathcal{P},\mathcal{R}$ (analogous to arbitrary low-complexity models $\Sigma$) and arbitrary distributions $\mathcal{A}$ for the forwards operator. Unsurprisingly, a number of the conditions in our main result, Theorem 1.1 – namely, the low-complexity condition (ii) and concentration condition (iii) – share similarities to those that ensure stable and accurate recovery in non-Bayesian compressed sensing. See Remarks 2.2 and 3 for further discussion. However, the proof techniques used in this work are once more entirely different.
Finally, while the focus of this work is to establish guarantees for posterior sampling, we mention in passing related work on information-theoretically optimal recovery methods. Methods such as Approximate Message Passing (AMP) donoho2009message ; bayati2011dynamics are well studied, and asymptotic information-theoretic bounds for Gaussian random matrices are known karan2024unrolled . AMP methods are fast point estimation algorithms, whereas we focus on sampling-based methods and do not consider computational implementations (see §5 for some further discussion). Note that information-theoretic lower bounds for posterior sampling have also been shown for the Gaussian case in jalal2021instance .